# Attention Mechanisms Don’t Learn Additive Models: Rethinking Feature Importance for Transformers

### v1

@@ 3.3 Encoder-Only and Decoder-Only models
Practical implementations introduce subtle modifications into the process described previously. The most relevant distinction is made between encoder-only models, that include BERT [15] and its variants, and decoder-only models such as the GPT models [38, 39].
Encoder-only models. Considering BERT as an example of an encoder-only model, the first token is used for the classification, i.e, $r=1$. Usually, a special token [CLS] is prepended to the text at position $1$, however this is not strictly necessary for the functioning of the model.
Decoder-only models. In contrast, decoder-models like GPT-2 [39] add the classification head on top of the last token for classification, i.e., $r={\lvert{\bm{t}}\rvert}$. A key difference is that in GPT-2 and other decoder-only models, a causal mask is laid over the attention matrix, resulting in $\alpha_{i,j}=0$ for $j>i$. This encodes the constraint that tokens can only attend to themselves or to previous ones.
@@ 4 Analysis
Note that at this point, we can let the sum go towards ${\lvert{\bm{t}}\rvert}$ for both encoder and decoder models. For encoder models, this is the default. For decoder models, we have ${\bm{s}}_{r}=\sum_{j=1}^{r}a_{rj}{\bm{v}}_{j}$ but as $r={\lvert{\bm{t}}\rvert}$, the outcome is the same as given in Eqn. 1.
@@ 4.1 Transformers cannot represent additive models
We now consider how this architecture would represent a linear model. In this model, each token is assigned a weight $w:\mathcal{V}\rightarrow\mathbb{R}$. The output obtained by adding weights and an offset $b\in\mathbb{R}$:
$F([t_{1},t_{2},\ldots,t_{\lvert{\bm{t}}\rvert}])\overset{!}{=}b+\sum_{i=1}^{{\lvert{\bm{t}}\rvert}}w(t_{i}).$ (2)
The transformer gives rise to an interesting observation when considering token sequences of identical tokens but of different lengths, i.e., $[\tau],[\tau,\tau],$ etc. We first note that the sum of the attention scores is bound to be $\sum_{j=1}^{{\lvert{\bm{t}}\rvert}}a_{rj}=1$. The output of the attention head will thus be a weighted average of the value vectors ${\bm{v}}_{i}$. As the value vectors ${\bm{v}}_{j}(t_{j})$ are determined purely by the input tokens $t_{j}$, for a sequence of identical tokens, we will have the same value vectors, resulting in identical vectors being averaged. In summary, this renders the transformer incapable to differentiate between sequences of different lengths. We are now ready to state our main result, which formalizes this intuition.
Proposition 4.1 (Single-layer transformers cannot represent generalized additive models (GAMs)).
Let $\mathcal{V}$ be a vocabulary and $C\geq 2,C\in\mathbb{N}$ be a maximum sequence length (context length). Let $w_{i}:\mathcal{V}\rightarrow\mathbb{R},\forall i\in 1,...,C$ be any mapping that assigns a token encountered at position $i$ a numerical score including at least one token $\tau\in\mathcal{V}$ with non-zero weight $w_{i}(\tau)\neq 0$ for some $i\in 2,\ldots,C$. Let $b\in\mathbb{R}$ be an arbitrary offset. Then, there exists no parametrization of the encoder or decoder single-layer transformer $F$ such that for every sequence ${\bm{t}}=[t_{1},t_{2},\ldots,t_{\lvert{\bm{t}}\rvert}]$ with length ${\lvert{\bm{t}}\rvert}\leq C$, the output of the transformer network is equivalent to $F([t_{1},t_{2},\ldots,t_{\lvert{\bm{t}}\rvert}])=b+\sum_{i=1}^{{\lvert{\bm{t}}\rvert}}w_{i}(t_{i})$.
@@
We prove the statement by concatenating the token $\tau$ to sequences of different length. We then show that the inputs to the final part $g$ of the transformer will be independent of the sequence length. Due to $g$ being deterministic, the output will also be independent of the sequence length. This is contradictory to the GAM model with a weight $w_{j}(\tau)\neq 0$ requiring different outputs for sequences of length $j{-}1$ and $j$. Formal proofs for all results can be found in Appendix B. ∎
In simple terms, the proposition states that the transformer cannot represent GAMs for sequences of more than one token apart from trivial constant functions. Importantly, the class of functions stated in the above theorem includes the prominent case of linear models in Eqn. 1, where each token has a certain weight $w$ independent of its position in the input vector (i.e., $w_{i}\equiv w,\forall i$, see Corollary B.2).
4.2 Transformer networks with multiple layers
@@
Corollary 4.2 (Multi-Layer transformers cannot learn linear models either).
Under the same conditions as in Proposition 4.1, a stack of multiple transformer blocks as in the model $F$ neither has a parametrization sufficient to represent the linear model.
Proof Sketch.
To see this, we first prove that for identical input embeddings ${\bm{H}}^{(l-1)}$, that do not change for varying sequence lengths (although their number changes), the outputs ${\bm{H}}^{(l)}$ will be identical as well and independent of the sequence length. Subsequently, we show that the final embeddings will also be independent of the sequence length, resulting in a contradiction as before. ∎
4.3 Practical Considerations
The above derivation contains two slight deviation from the transformer architecture deployed in practice, that are of of minor importance for the validity of our results. First, common models such as BERT use a special token referred to as CLS-token where the classification head is placed on. In the example presented in the proof, the sequences have to be extended by this token as in ${\bm{t}}_{A}=[t_{\text{CLS}},\tau]$ and ${\bm{t}}_{B}=[t_{\text{CLS}},\tau,\tau]$. Thus, a small part of the attention will reside on this token as well.
To represent token order, common architectures use positional embeddings, tying the embedding vectors to the token position $i$. The behavior that we show in this work’s analysis does however also govern transformers with positional embeddings for the following reason: While the positional embeddings could be used by the non-linear ffn part to differentiate sequences of different length in theory, our proofs show that to represent the linear model, the softmax operation must be inverted for any input sequence. This is highly nonlinear operation and the number of possible sequences grows exponentially at a rate of $\lvert\mathcal{V}\rvert^{C}$ with the context length $C$. Learning-theoretic considerations (e.g., 6) show that the number of input-output pairs the two-layer networks deployed can maximally represent is bounded by $\mathcal{O}\left(dn_{\text{hidden}}\log(dn_{\text{hidden}})\right)$, which is small ($d{=}786,n_{\text{hidden}}{=}3072$ for BERT) in contrast to the number of sequences ($C=1024$, $\lvert\mathcal{V}\rvert\approx 3\times 10^{4}$). We conclude that the inversion is therefore impossible for realistic setups and positional embeddings can be neglected, which is confirmed by our empirical findings.
@@ 5 A Surrogate Model for Transformers
In the previous section, we theoretically established that transformer models struggle to represent additive functions. Note that this must not necessarily be considered a weakness. However, it certainly casts doubts on the suitability of additive models as surrogate models for explanations of transformers. For a principled approach, we consider the following four requirements to be of importance:
@@
    Interpretability. The surrogate model’s parameters should be inherently interpretable.
@@
    Learnability. The surrogate model should be easiliy representable by common transformers.
@@
    Recovery. If the predictive model falls into the surrogate model’s class, the fitted surrogate model’s parameters should match those of the predictive model.
@@
Figure 3: Transformers fail to learn linear models. We train different models on a synthetically sampled dataset where the log-odds obey a linear relation to the features. Fully connected models (2-layer ReLU networks with different hidden layer widths) capture the linear form of the relationship well despite some estimation error (a). However, common transformer models fail to model this relationship and output almost constant values (b)-(d). This does not change with more layers.
5.1 The Softmax-Linked Additive Log-Odds Model
To meet the requirements, we propose a novel discriminative surrogate model that explicitly models the behavior of the softmax function. Instead of only assigning a single weight $w$ to each token, we separate two characteristics: We introduce the token importance as a mapping $s:\mathcal{V}\rightarrow\mathbb{R}$ and a token value in form of a mapping $v:\mathcal{V}\rightarrow\mathbb{R}$. Subsequently, we consider the following discriminative model:
@@ 5.1 The Softmax-Linked Additive Log-Odds Model
Due to the shift invariance of the softmax function, we observe that the mappings $s$ and $s^{\prime}$ given by $s^{\prime}(\tau)=s(\tau)+\delta$ result in the same softmax-score and thus the same log-odds model for any input ${\bm{t}}$. Therefore, the parameterization would not be unique. To this end, we introduce a normalization constraint on the sum of token importances for uniqueness. Formally, we constrain it to a user-defined constant $\gamma\in\mathbb{R}$ such that $\sum_{\tau\in\mathcal{V}}s(\tau)=\gamma$, where natural choices include $\gamma\in\{{0,1\}}$. We refer to the discriminative model given in Eqn. (4) together with the normalization constraint as the softmax-linked additive log-odds model (SLALOM).
@@ 5.2 Theoretical properties of SLALOM
For any mapping $s$, $v$ and a transformer with an embedding size $d$ and head dimension $d_{h}$ with $d,d_{h}\geq 3$, there exists a parameterization of the transformer to reflect SLALOM in Equation 4 together with the normalization constraint.
Proof Sketch.
We prove this statement by explicitly constructing the weight matrices of the transformer for this mapping.In particular, this involves obtaining raw attention scores of $s(t_{j})$ for key $t_{j}$ that are normalized as in the SLALOM model. See the full proof for the full construction in the appendix. ∎
This proposition highlights that – unlike linear models – there are simple ways for the transformer to represent relations governed by SLALOMs. We demonstrate this empirically in our experimental section and conclude that SLALOM fulfills Property (2). For Property (3), Recovery, we make the following proposition:
@@
Suppose query access to a model $G$ that takes sequences of tokens ${\bm{t}}$ with lengths $\lvert{\bm{t}}\rvert\in 1,\ldots,C$ and returns the log-odds according to a non-constant SLALOM on a vocabulary $\mathcal{V}$, normalization constant $\tau\in\mathbb{R}$, but with unknown parameter mappings $s:\mathcal{V}\rightarrow\mathbb{R}$, $v:\mathcal{V}\rightarrow\mathbb{R}$. For $C\geq 2$, we can recover the true mappings $s$, $v$ with $12\lvert\mathcal{V}\rvert-$ 1 forward passes of $F$.
Proof Sketch.
We first query $G([\tau]),\forall\tau\in\mathcal{V}$ for single token sequences. We know that for ${\lvert{\bm{t}}\rvert}=1$ all attention is on one token, i.e., $(\alpha_{i}=1)$ and we thus have $G([\tau])=v(\tau)$. Having the values of each token, We then query $n-1$ two-token sequences with $G([\tau,\omega]),\omega\in\mathcal{V}\setminus\{\tau\}$. Given the token’s value scores, we can obtain their relative importances using $n-1$ forward passes. We use the condition that the SLALOM is non-constant to make sure $v(\tau)$ and $v(\omega)$ are always different and we can use the output to obtain the importances mapping $s$. ∎
This statement confirms property (3) and shows that SLALOM can be uniquely reidentified when we rule out the corner case of constant models.
Complexity considerations. Computational complexity can be a concern for XAI methods. To estimate exact Shapley values, the model’s output on exponentially many feature coalitions needs to be evaluated. However, as the proof of Proposition 5.2 shows, to estimate SLALOM’s parameters for an input sequence of $\mathcal{V}$ tokens, only $12\lvert\mathcal{V}\rvert-$ 1 forward passes are required, verifying Property (4). We empirically show that computing SLALOM explanations is about 8$\times$ faster than computing LIME or SHAP explanations using the same number of samples (cf. Section F.7 for results and details).
@@ 5.3 Numerical Algorithms for computing SLALOMs
We suggest two algorithms to fit SLALOMs post-hoc on input-output pairs of a trained model:
SLALOM-Ana. We can use the analytical solution given in the proof of Proposition 5.2. However, the algorithm is numerically unstable as it uses a small number of short sequences (cf. Section F.3).
SLALOM-SGD. To this end we use SGD to optimize SLALOM’s parameters such that the logit output of the transformer closely resembles that of the surrogate SLALOM on a number of query samples. To explain a specific sample, we obtained good results when fitting SLALOM on random sequences made up of its tokens (see Appendix D for pseudocode).
(a) BERT
(b) GPT-2
(c) Recovering Log-Odds
(d) Recovering Params.
Figure 4: SLALOM describes outputs of transformer models well (a, b). Fitting SLALOM to the outputs of the models shown in Figure 3. Despite having $C/2{=}15\times$ more parameters than the SLALOM model, the GAM model does not match the transformer behavior. We provide quantitative results in Table 6 (Appendix). Recovering SLALOMs. We observe that we can recover the original logit-scores by fitting SLALOM on a 2-layer DistilBERT model (c) and we see a strong connection between original SLALOM parameters and the recovered ones (d). More results in Section F.2.

### v2

@@ 3.3 Encoder-only and decoder-only models
Practical architectures can be seen as parametrizations of the process described previously. We introduce the ones relevant to this work in this section. Commonly, a distinction is made between encoder-only models, that include BERT (Devlin et al., 2018) and its variants, and decoder-only models such as the GPT models (Radford et al., 2018; 2019).
Encoder-only models. Considering BERT as an example of an encoder-only model, the first token is used for the classification head, i.e., $r=1$. Usually, a special token [CLS] is prepended to the text at position $1$. However this is not strictly necessary for the functioning of the model.
Decoder-only models. In contrast, decoder-only models like GPT-2 (Radford et al., 2019) add the classification head on top of the last token for classification, i.e., $r={\lvert{\bm{t}}\rvert}$. A key difference is that in GPT-2 and other decoder-only models, a causal mask is laid over the attention matrix, resulting in $\alpha_{i,j}=0$ for $j>i$. This encodes the constraint that tokens can only attend to themselves or to previous ones.
To make our model amenable to theoretical analysis, the transformer model in our analysis contains one slight deviation from practical models. We do not consider positional embeddings, which are added to the embeddings based on their position in the sentence. We will empirically demonstrate that using positional embeddings does not affect the validity of our findings on practical models.
@@ 4 Analysis
The attention output is given by a sum of the token value vectors ${\bm{v}}_{j}$ weighted by the respective attention weights $\alpha_{rj}$.
@@ 4.1 Transformers cannot represent additive models
We now consider how this architecture would represent a linear model. In this model, each token is assigned a weight $w:\mathcal{V}\rightarrow\mathbb{R}$. The output is obtained by adding weights and an offset $b\in\mathbb{R}$, consequently requiring
$F([t_{1},t_{2},\ldots,t_{\lvert{\bm{t}}\rvert}])=b+\sum_{i=1}^{{\lvert{\bm{t}}\rvert}}w(t_{i})$ (2)
for all possible input sequences.
The transformer shows surprising behavior when considering sequences of identical tokens but of different lengths, i.e., $[\tau],[\tau,\tau],\ldots$ We first note that the sum of the attention scores is bound to be $\sum_{j=1}^{{\lvert{\bm{t}}\rvert}}a_{rj}=1$. The output of the attention head will thus be a weighted average of the value vectors ${\bm{v}}_{i}$. We find that the first-layer value vectors ${\bm{v}}_{j}(t_{j})$ are determined purely by the input tokens $t_{j}$ (cf. full formalization in Section B.1). For a sequence of identical tokens, we will thus have the same value vectors, resulting in identical vectors being averaged. This makes the transformer produce the same output for each of these sequences. This contradicts the form in (2), where the output should successively increase by $w(t_{i})$. We are now ready to state our result, which formalizes this intuition for the more general class of additive models where the token weight may also depend on its position $i$ in the input (equivalent to a GAM).
Proposition 4.1 (Single-layer transformers cannot represent additive models.).
Let $\mathcal{V}$ be a vocabulary and $C\geq 2,C\in\mathbb{N}$ be a maximum sequence length (context length). Let $w_{i}:\mathcal{V}\rightarrow\mathbb{R},\forall i\in 1,...,C$ be any map that assigns a token encountered at position $i$ a numerical score including at least one token $\tau\in\mathcal{V}$ with non-zero weight $w_{i}(\tau)\neq 0$ for some $i\in 2,\ldots,C$. Let $b\in\mathbb{R}$ be an arbitrary offset. Then, there exists no parametrization of the encoder or decoder single-layer transformer $F$ such that for every sequence ${\bm{t}}=[t_{1},t_{2},\ldots,t_{\lvert{\bm{t}}\rvert}]$ with length ${\lvert{\bm{t}}\rvert}\leq C$, the output of the transformer network is equivalent to $F([t_{1},t_{2},\ldots,t_{\lvert{\bm{t}}\rvert}])=b+\sum_{i=1}^{{\lvert{\bm{t}}\rvert}}w_{i}(t_{i})$.
@@
We prove the statement by concatenating the token $\tau$ to sequences of different length. We then show that the inputs to the final part $g$ of the transformer will be independent of the sequence length. Due to $g$ being deterministic, the output will also be independent of the sequence length. This is contradictory to the additive model with a weight $w_{j}(\tau)\neq 0$ requiring different outputs for sequences of length $j{-}1$ and $j$. Formal proofs for all results can be found in Appendix B. ∎
In simple terms, the proposition states that the transformer cannot represent any additive models on sequences of more than one token besides constant functions or those fully determined by the first input token. Importantly, the class of functions stated in the above theorem includes the prominent case of linear models in Eqn. 1, where each token has a certain weight $w$ independent of its position in the input vector (i.e., $w_{i}\equiv w,\forall i$, see Corollary B.2). We would like to emphasize that this statement includes the converse:
Corollary 4.2.
Transformers whose outputs are not constant or fully determined by the first token of the input sequence cannot be functionally equivalent to an additive model.
4.2 Transformer networks with multiple layers cannot represent additive models
@@
Corollary 4.3 (Multi-Layer transformers cannot learn additive models either).
Under the same conditions as in Proposition 4.1, a stack of multiple transformer blocks as the model $F$, neither has a parametrization sufficient to represent the additive model.
Practical considerations. As stated earlier, the transformer model in our analysis does not consider positional embeddings that are added on the token embeddings. However, this does not have major ramifications in practice: While the transformer would be able to differentiate between sequences of different lengths with positional embeddings in theory, the softmax operation must be inverted for any input sequence by the linear feed-forward block that follows the attention mechanism. This is a highly non-linear operation and the number of possible sequences grows exponentially with the context length and vocabulary size. Learning-theoretic considerations suggest that this inversion is impossible for reasonably-sized networks as outlined in Section C.2. We will confirm our results with empirical findings obtained exclusively on non-modified models with positional embeddings.
@@ 5 A Surrogate Model for Transformers
In the previous section, we theoretically established that transformer models struggle to represent additive functions. While this must not necessarily be considered a weakness, it certainly casts doubts on the suitability of additive models as surrogate models for explanations of transformers. For a principled approach, we consider the following four requirements to be of importance:
@@
    Interpretability. The surrogate model should be simple enough such that its parameters are inherently interpretable for humans (Molnar, 2019, Chapter 9.2).
@@
    Learnability. The surrogate model should be easily representable by common transformers. Learnability is crucial because using a surrogate model that is hard to represent for the predictive model will likely result in low-fidelity explanations.
@@
    Recovery. If the predictive model falls into the surrogate model’s class, the fitted surrogate model’s parameters should match those of the predictive model. Together with learnability, recovery ensures that a model can pick up the true relations present in the data and they can be re-identified by the explanation, which is essential, e.g., in scientific discovery with XAI.
@@
Figure 3: Transformers fail to learn linear models. We train different models on a synthetically sampled dataset where the log odds obey a linear relation to the features. Fully connected models (2-layer ReLU networks with different hidden layer widths) capture the linear form of the relationship well despite some estimation error (a). However, common transformer models fail to model this relationship and output almost constant values (b)-(d). This does not change with more layers.
5.1 The Softmax-Linked Additive Log Odds Model
To meet the requirements, we propose a novel discriminative surrogate model that explicitly models the behavior of the softmax function. Instead of only assigning a single weight $w$ to each token, we separate two characteristics: We introduce the token importance as a map $s:\mathcal{V}\rightarrow\mathbb{R}$ and a token value in form of a map $v:\mathcal{V}\rightarrow\mathbb{R}$. Subsequently, we consider the following discriminative model:
@@ 5.1 The Softmax-Linked Additive Log Odds Model
Due to the shift-invariance of the softmax function, we observe that the maps $s$ and $s^{\prime}$ given by $s^{\prime}(\tau)=s(\tau)+\delta$ result in the same softmax-score and thus the same log odds model for any input ${\bm{t}}$. Therefore, the parameterization would not be unique. To this end, we introduce a normalization constraint on the sum of token importances for uniqueness. Formally, we constrain it to a user-defined constant $\gamma\in\mathbb{R}$ such that $\sum_{\tau\in\mathcal{V}}s(\tau)=\gamma$, where natural choices include $\gamma\in\{{0,1\}}$. We refer to the discriminative model given in Eqn. (4) together with the normalization constraint as the softmax-linked additive log odds model (SLALOM).
@@ 5.2 Theoretical properties of SLALOM
For any maps $s$, $v$, and a transformer with an embedding size $d$ and head dimension $d_{h}$ with $d,d_{h}\geq 3$, there exists a parameterization of the transformer to reflect SLALOM in Equation 4 together with the normalization constraint.
This statement can be proven by explicitly constructing the corresponding weight matrix (cf. Section B.5). This proposition highlights that – unlike linear models – there are simple ways for the transformer to represent relations governed by SLALOMs. We demonstrate this empirically in our experimental section and conclude that SLALOM fulfills Property (2). For Property (3), Recovery, we make the following proposition:
@@
Suppose query access to a model $G$ that takes sequences of tokens ${\bm{t}}$ with lengths $\lvert{\bm{t}}\rvert\in 1,\ldots,C$ and returns the log odds according to a non-constant SLALOM on a vocabulary $\mathcal{V}$, normalization constant $\tau\in\mathbb{R}$, but with unknown parameter maps $s:\mathcal{V}\rightarrow\mathbb{R}$, $v:\mathcal{V}\rightarrow\mathbb{R}$. For $C\geq 2$, we can recover the true maps $s$, $v$ with $12\lvert\mathcal{V}\rvert-$ 1 forward passes of $F$.
This statement confirms property (3) and shows that SLALOM can be uniquely re-identified when we rule out the corner case of constant models. We prove it in Section B.6.
Complexity considerations. Computational complexity can be a concern for XAI methods. To estimate exact Shapley values, the model’s output on exponentially many feature coalitions needs to be evaluated. However, as the proof of Proposition 5.2 shows, to estimate SLALOM’s parameters for an input sequence of $\mathcal{V}$ tokens, only $12\lvert\mathcal{V}\rvert-$ 1 forward passes are required, verifying Property (4). We empirically show that computing SLALOM explanations is about 5$\times$ faster than computing SHAP explanations when using the same number of samples in our experimental section.
@@ 5.3 Numerical algorithms for computing SLALOMs
Having derived SLALOM as a better surrogate model, we require numerical algorithms to estimate $v$ and $s$. Unfortunately, the strategy derived in Proposition 5.2 using a minimal number of samples is numerically unstable. We make two key implementation choices for SLALOM to be used as an explanation technique. First, we can control the sample set of features and labels obtained through queries of the predictive model. Second, we can use different optimization strategies to fit SLALOM on this sample set. We suggest two algorithms to fit SLALOMs post-hoc on input-output pairs of a trained predictive model:
SLALOM-eff. The first version of the algorithm to fit SLALOM models is designed for maximum efficiency while maintaining reasonable performance across several XAI metrics. Obtaining a large dataset of input-output pairs can incur substantial computational costs as a forward pass of the models needs to be executed for each sample. To speed up this process, SLALOM-Eff uses very short sequences (we use only two tokens in this work) randomly sampled from the vocabulary for this purpose. To efficiently fit the surrogate model, we perform stochastic gradient descent on SLALOM’s parameters using the mean-squared-error loss between the score output by SLALOM and the score by the predictive model as an objective. SLALOM-eff is our default technique used unless stated otherwise.
SLALOM-fidel. We provide another technique to fit SLALOM optimized for maximum fidelity under input perturbations such as token removals. To explain a specific sample, we sample input where we remove up to $K$ randomly sampled tokens. The sequences with tokens removed and their predictive model scores are used to fit the model, similar to LIME (Ribeiro et al., 2016). Instead of SGD, we can leverage optimizers for Least-Square-Problems to fit the parameters iteratively, however incurring a higher latency. We provide details and pseudocode for both fitting routines in Appendix D.
(a) Fitting SLALOM to BERT outputs
(b) Fitting SLALOM to GPT-2 outputs
(c) Comparing SLALOM and predictive output
(d) Recovering parameters of SLALOM dataset
Figure 4: Verifying properties with synthetic data: SLALOM describes outputs of transformer models well (a, b). We fit SLALOM to the outputs of the BERT and GPT-2 models trained on the linear synthetic dataset. The linear and GAM models (despite having $C/2{=}15\times$ more parameters) do not match the transformer’s behavior. We provide another empirical counterexample and additional quantitative results in Section F.1. Verifying recovery (c, d). We verify the recovery property on a second synthetic dataset where features and labels obey a SLALOM relation. We train a 2-layer DistilBERT model on the data and fit SLALOM to the trained model. We can recover the original logit scores (c) and see a strong connection between original SLALOM parameters and the recovered ones (d). These findings verify the learnability and recovery properties. More results in Section F.2.